# SeedProteo: Accurate De Novo All-Atom Design of Protein Binders

### v1

@@ Abstract
We present SeedProteo, a diffusion-based model for de novo all-atom protein design. We demonstrate how to repurpose a cutting-edge folding architecture into a powerful generative design framework by effectively integrating self-conditioning features. Extensive benchmarks highlight the model’s capabilities across two distinct tasks: in unconditional generation, SeedProteo exhibits superior length generalization and structural diversity, maintaining robustness for long sequences and complex topologies; in binder design, it achieves state-of-the-art performance among open-source methods, attaining the highest in-silico design success rates, structural diversity and novelty.
@@
Figure 1: Comparison of binder design success and diversity across ten target proteins. We benchmark exclusively against open-source methods using SeedFold in-silico metrics.
@@ 1 Introduction
In this study, we introduce SeedProteo, a diffusion-based model that learns the manifold of proteins at the all-atom level. SeedProteo mimics the architecture of AlphaFold3 [abramsonAccurateStructurePrediction2024, wohlwendBoltz1DemocratizingBiomolecular2024, passaroAccurateEfficientBinding], which has proven its effectiveness in protein folding. Through unifying the representation of different amino acids, we significantly enhance its generative power when the protein sequence and homology are not provided. A major issue in all-atom modeling is the backbone-sidechain inconsistency (i.e., the sequence-structure inconsistency), which means that the sidechain atoms are locally plausible but the derived sequence fails to fold into the global backbone structure. SeedProteo iteratively evolves the sequence and structure and integrates two components to enhance the modeling ability of two modalities: (i) SeedProteo enhances the structural feasibility through the integration of geometric reasoning in the denoising process. To be specific, we reuse the denoised structure from the previous timestep [chen2022analog] to stabilize the sampling process and conduct geometric reasoning in the pair representation space. (ii) SeedProteo conducts energy minimization within the sequence space modeled by a Markov random field (MRF) [metzler2005markov]. By integrating high-order couplings among distinct positions, we are able to sample globally feasible sequences and prevent being trapped in the local minima of simple patterns constituted by high-frequency amino acids. We conduct experiments on two tasks: unconditional sampling and binder design. The results demonstrate that:
@@ 2.2 Self-conditioning features
    Sequence Sampling: SeedProteo employs a Markov Random Field (MRF)-based Sequence Module [renAccurateRobustProtein2024] to decode amino acid sequences. The probability of a sequence $\mathbf{x}$ is parameterized using the learned pair and single representations (Eq. 1):
@@ 3.2 Binder Design Benchmark
Beyond design success, structural diversity is a critical metric for downstream applicability. Figure 1 highlights that SeedProteo-D excels in generating structurally distinct binders, achieving the highest number of unique success clusters in 8 out of the 10 targets. Complementing this structural diversity, Table 2 demonstrates that SeedProteo-D also achieves the best novelty scores for successfully designed binders in 8 out of the 10 targets. Notably, it consistently outperforms all baselines, except RFDiffusion3, across every target; for example, while baselines like PXDesign often yield scores above 0.9, SeedProteo-D maintains superior novelty with scores in the 0.8 range. This indicates that SeedProteo does not merely converge on a single solution but effectively explores the structural landscape to provide a diverse portfolio of high-quality and novel candidates. Figure 4 presents case studies on challenging targets.
@@ 3.3 Discussion on Co-Design Strategies
Empirical results (Figure 1 and Table 3) indicate that, on average, sequences derived from ProteinMPNN achieve higher success rates than those generated via pure co-design. This is expected, as ProteinMPNN is highly optimized for fixed-backbone recovery. However, we argue that reliance on inverse folding alone is not the ultimate solution for de novo binder design.
@@ 4 Conclusion
In this work, we presented SeedProteo, a diffusion-based all-atom protein design model. Our empirical benchmarks demonstrate that SeedProteo exhibits superior robustness across long sequences and complex topologies in unconditional design. Furthermore, it achieves state-of-the-art performance among open-source methods for binder design, particularly on hard and multi-chain targets. While effective co-design remains an open challenge in the field, we offer specific insights into achieving superior sequence-structure consistency via a global MRF module. We hope these findings will inspire significant community efforts to further advance the frontier of simultaneous sequence-structure generation.
@@ Acknowledgments
[…]
5 Data Preparation and Curation
Training a foundation model for protein design, particularly one capable of binder generation, necessitates a diverse corpus comprising both monomeric and multimeric structures. While predicted monomer databases such as the AlphaFold Database (AFDB) [varadiAlphaFoldProteinStructure2022a] and ESMAtlas [lin2023evolutionary] provide extensive coverage of single-chain structures, existing multimer datasets derived from the RCSB PDB [berman2000protein] are limited in scale and often unsuitable for the specific two-chain requirements of binder design tasks. To address this, we constructed a composite training dataset by integrating curated monomer data with multimeric interactions sourced from Pinder [kovtun2024pinder] and HumanPPI [zhangPredictingProteinproteinInteractions2025]. We detail the processing and filtering pipelines for each category below.
5.1 Monomer Data Curation
Our monomeric dataset is derived exclusively from AFDB and ESMAtlas. For the ESMAtlas component, to mitigate potential quality issues associated with ESMFold [lin2023evolutionary] predictions, we retrieved the corresponding sequences from the MGnify database [richardson2023mgnify] and re-predicted their structures using AlphaFold2 (AF2) [jumper2021highly].
To ensure high data quality, we applied the following inclusion criteria to all monomer structures:
  • Sequence length between 50 and 768 residues.
  • Average pLDDT score >80>80.
  • Secondary structure coil fraction $50\%$.
To eliminate redundancy, we performed clustering using Foldseek [vankempenFastAccurateProtein2024] and MMseqs2 [steineggerMMseqs2EnablesSensitive2017] and retained only the cluster representatives (centroids). The final curated monomer dataset consists of approximately 0.5 million structures. Sequences that can be folded by SeedFold [seedfold] in a single-sequence setting are exclusively allocated for the training of the initial phase.
5.2 Multimer Data Curation
Protein-Protein Interactions (PPIs) are fundamental to training a robust binder design model. We sourced multimer data from two primary streams to cover experimentally determined interactions and high-quality predicted domain interactions.
Experimental PPIs (Pinder).
To align our training data with the standard inference scenario, where a binder chain is designed for a specific target chain, we utilized the Pinder dataset. Using the curated holo-structure PDB IDs and chain IDs provided by Pinder, we extracted relevant substructures from RCSB biological assemblies. We filtered these two-chain complexes by requiring: (1) a coil fraction of $50\%$ for each chain, and (2) a minimum interfacial Cβ distance of $8\text{\AA }$ to ensure meaningful contact. Following structural clustering via Foldseek-Multimer, we retained approximately 50,000 unique cluster representatives.
Augmented DDI Dataset (HumanPPI).
To overcome the scarcity of experimental PPI data, we augmented our training set with domain-domain interactions (DDIs). This approach relies on the established hypothesis that intrachain DDI interfaces resemble interchain PPI interfaces in terms of coevolutionary and physicochemical properties [lau2024exploring]. We utilized the HumanPPI database, which distills DDIs from accurately predicted monomer structures in the AFDB. We applied the identical filtering and clustering pipeline used for the Pinder dataset (coil check and distance constraints). This augmentation yielded an additional $\sim 0.1$ million interaction pairs.
Table 5: Training Stages.
Hyperparameter Stage 1 (Initial) Stage 2 (Fine-tuning 1) Stage 3 (Fine-tuning 2)
Crop Size 384 768 768
Batch Size 128 64 64
Motif Percentage 0% 20% 20%
Training Steps 50K 20K 30K
Data Sampling Distribution
    Strictly Filtered Monomer 100% 20% 10%
    Expanded Monomer Data 80% 40%
    Multimer Data 50%
6 Evaluation Settings
6.1 Unconditional Generation Evaluation
[…]
Topology-Aware Analysis.
To address the long-standing challenge in protein generative models, specifically the bias toward $\alpha$-helical structures and the difficulty in generating valid $\beta$-sheets, we incorporate a fine-grained analysis based on secondary structure topology. Based on secondary structure proportions, we classify the generated proteins into three distinct fold types. We define EEE (primarily $\beta$-sheet) as structures with over 40% sheet and under 20% helix content, and HHH (primarily $\alpha$-helical) as those with over 50% helix and under 10% sheet content; both categories require a loop fraction below 45%. All remaining structures are categorized as hybrid Helix-E-Loop (HEL).
6.2 Binder Design Evaluation
We evaluate binder design performance on a benchmark set of 10 targets, consistent with the validation set used in AlphaProteo (see Table 6). To rigorously assess the upper bound of design capability across different length scales, we adopt a dense sampling strategy: for each target, we sample 100 candidates at every 5-residue interval within the target-specific binder length range. For example, for the SC2RBD target (length range 80–120 residues), this results in a total of 900 sampled candidates.
For sequence recovery, we generate two sequences for each designed binder backbone using SolubleMPNN at a sampling temperature of $\tau=0.001$. The designs are then validated using SeedFold in single sequence with target template mode. A binder design is considered successful if it meets the following criteria:
  • Interface Predicted Aligned Error (min PAE interaction) <1.5<1.5;
  • Binder pTM score >0.8>0.8;
  • Complex RMSD <2.5<2.5Å.
Table 6: Benchmark Dataset for Binder Design. We evaluated performance on 10 diverse targets. The table lists the target details, including the PDB ID, the specific chains/residues used as the target region, the key hotspot residues defining the binding site, and the range of binder lengths sampled during inference.
Target PDB ID Target Region Hotspot Residues Length Range
BHRF1 2wh6 Chain A: 2-158 A65, A74, A77, A82, A85, A93 80-120
SC2RBD 6m0j Chain E: 333-526 E485, E489, E494, E500, E505 80-120
IL-7RA 3di3 Chain B: 17-209 B58, B80, B139 50-120
PD-L1 5o45 Chain A: 17-132 A56, A115, A123 50-120
TrkA 1www Chain X: 282-382 X294, X296, X333 50-120
Insulin 4zxb Chain E: 6-155 E64, E88, E96 40-120
H1 5vli Chain A: 1-50, 76-80, 107-111, 258-322; Chain B: 1-68, 80-170 B21, B45, B52 40-120
VEGF-A 1bj1 Chain V: 14-107; Chain W: 14-107 W81, W83, W91 50-140
IL-17A 4hsa Chain A: 17-131; Chain B: 19-127 A94, A116, B67 50-140
[…]
7 Baseline Sampling
In this section, we detail the specific configurations, checkpoints, and codebases used for all baseline methods included in our benchmarks.
7.1 Unconditional Benchmark
La-Proteina. We utilized the official implementation and checkpoints provided in the La-Proteina repository [^1] https://github.com/NVIDIA-Digital-Bio/la-proteina. To ensure optimal performance across different scales, we selected checkpoints based on the target sequence length: the LD1_ucond_notri_512.ckpt checkpoint was used for lengths 100–500, while LD3_ucond_notri_800.ckpt was employed for lengths 600–1000. All sampling was conducted using the default inference configurations.
Proteina. We used the code and checkpoints from the official Proteina repository [^2] https://github.com/NVIDIA-Digital-Bio/proteina. Similar to La-Proteina, we adopted a length-dependent checkpoint strategy: proteina_v1.1_DFS_200M_tir.ckpt was used for lengths 100–500, and proteina_v1.6_DFS_200M_notri_long_chain_generation.ckpt was used for lengths 600–1000. Default sampling parameters were applied in all cases.
RFDiffusion. We employed the official codebase available at the RFDiffusion repository [^3] https://github.com/RosettaCommons/RFdiffusion. All unconditional generation tasks were performed using the standard default sampling configurations provided by the authors.
7.2 Binder Design Benchmark
For the binder design task, all methods utilize identical target structures (in mmCIF format) and are conditioned on the exact same set of hotspot residues. For each target, we generate an equal number of candidate proteins with consistent lengths across all models. To ensure a fair comparison of generative capabilities, we disable any built-in auxiliary modules—such as pre-filtering, ranking, or early-stopping strategies—thereby restricting the evaluation strictly to the core sampling function.
BindCraft. We utilized the official implementation from the BindCraft repository [^4] https://github.com/martinpacesa/BindCraft. For binder generation, we adopted the default_4stage_multimer configuration under the advanced setting. To ensure a fair evaluation of the generative capability without post-hoc selection bias, we disabled both the filtering stage and the early stopping strategy. This ensured that every sampling attempt was completed and all raw outputs were retained, regardless of their internal quality metrics.
BoltzGen. We performed evaluations using the code provided in the BoltzGen repository [^5] https://github.com/HannesStark/boltzgen, specifically utilizing commit SHA 58c1eed. We employed the protein-anything protocol for inference. To bypass internal ranking and filtering mechanisms, we set the inference budget equal to the number of desired designs, ensuring a direct one-to-one output.
RFDiffusion. For binder design tasks, we set both noise_scale_ca and noise_scale_frame to 0 during sampling. The authors suggest that this deterministic setting yields superior performance for specific binder design scenarios.
RFDiffusion3. We utilized the implementation available in the RosettaCommons Foundry repository [^6] https://github.com/RosettaCommons/foundry, specifically commit SHA 7f6656e. We set align_trajectory_structures=False to prevent potential runtime errors during trajectory processing; all other configurations were maintained at default values.
PXDesign. We utilized the implementation in the PXDesign repository [^7] https://github.com/bytedance/PXDesign, specifically commit SHA ec6615c. We perform binder design for each target of a given hotspot using Generation Only mode.

### v2

@@ Abstract
We present SeedProteo, a diffusion-based model for de novo all-atom protein design. We demonstrate how to repurpose a cutting-edge folding architecture into a powerful generative design framework by effectively integrating self-conditioning features. Extensive benchmarks highlight the model’s capabilities across two distinct tasks: in unconditional generation, SeedProteo exhibits superior length generalization and structural diversity, maintaining robustness for long sequences and complex topologies; in binder design, it achieves state-of-the-art performance among open-source methods, attaining the highest in-silico design success rates, structural diversity and novelty. Finally, we validate SeedProteo through wet-lab assays on two therapeutic targets, achieving hit rates of 70%–80% and picomolar-level binding affinities, establishing leading results. To facilitate community adoption, we provide public access to SeedProteo via a webserver (https://seedfold.io/proteinDesign).
@@ 1 Introduction
In this study, we introduce SeedProteo, a diffusion-based model that learns the manifold of proteins at the all-atom level. SeedProteo mimics the architecture of AlphaFold3 [abramsonAccurateStructurePrediction2024, wohlwendBoltz1DemocratizingBiomolecular2024, passaroAccurateEfficientBinding], which has proven its effectiveness in protein folding. Through unifying the representation of different amino acids, we significantly enhance its generative power when the protein sequence and homology are not provided. A major issue in all-atom modeling is the backbone-sidechain inconsistency (i.e., the sequence-structure inconsistency), which means that the sidechain atoms are locally plausible but the derived sequence fails to fold into the global backbone structure. SeedProteo iteratively evolves the sequence and structure and integrates two components to enhance the modeling ability of two modalities: (i) SeedProteo enhances the structural feasibility through the integration of geometric reasoning in the denoising process. To be specific, we reuse the denoised structure from the previous timestep [chen2022analog] to stabilize the sampling process and conduct geometric reasoning in the pair representation space. (ii) SeedProteo conducts energy minimization within the sequence space modeled by a Markov random field (MRF) [metzler2005markov]. By integrating high-order couplings among distinct positions, we are able to sample globally feasible sequences and prevent being trapped in the local minima of simple patterns constituted by high-frequency amino acids. We conduct extensive experiments on unconditional sampling, binder design, and in vitro validation. The results demonstrate that:
@@
4) Beyond in-silico evaluation, we further validate SeedProteo with wet-lab experiments, demonstrating that designed binders exhibit strong binding in SPR assays with high hit rates across multiple therapeutic targets.
@@ 2.2 Self-conditioning features
    Sequence Sampling: SeedProteo employs a Markov Random Field (MRF)-based Sequence Module [renAccurateRobustProtein2024, meynardfast] to decode amino acid sequences. The probability of a sequence $\mathbf{x}$ is parameterized using the learned pair and single representations (Eq. 1):
@@ 3.2 Binder Design Benchmark
Beyond design success, structural diversity is a critical metric for downstream applicability. Figure SeedProteo: Accurate De Novo All-Atom Design of Protein Binders highlights that SeedProteo-D excels in generating structurally distinct binders, achieving the highest number of unique success clusters in 8 out of the 10 targets. Complementing this structural diversity, Table 2 demonstrates that SeedProteo-D also achieves the best novelty scores for successfully designed binders in 8 out of the 10 targets. Notably, it consistently outperforms all baselines, except RFDiffusion3, across every target; for example, while baselines like PXDesign often yield scores above 0.9, SeedProteo-D maintains superior novelty with scores in the 0.8 range. This indicates that SeedProteo does not merely converge on a single solution but effectively explores the structural landscape to provide a diverse portfolio of high-quality and novel candidates. Figure 3 presents case studies on challenging targets.
@@ 3.3 Discussion on Co-Design Strategies
Empirical results (Figure SeedProteo: Accurate De Novo All-Atom Design of Protein Binders and Table 3) indicate that, on average, sequences derived from ProteinMPNN achieve higher success rates than those generated via pure co-design. This is expected, as ProteinMPNN is highly optimized for fixed-backbone recovery. However, we argue that reliance on inverse folding alone is not the ultimate solution for de novo binder design.
@@ 3.4 Ablation Study
4 In Vitro Experiments
Figure 4: Wet-lab validation of SeedProteo-designed binders. For each target (PD-L1, top; SC2RBD, bottom), three representative binding complex structures are shown alongside the SPR sensorgram of the highest-affinity binder.
To validate the practical utility of SeedProteo beyond in-silico metrics, we conducted wet-lab experiments on two therapeutically relevant protein targets: the SARS-CoV-2 receptor-binding domain (SC2RBD) [cao2020novo] and Programmed Death-Ligand 1 (PD-L1) [yang2025design]. These targets were selected due to their high clinical relevance—SC2RBD is a primary target for COVID-19 therapeutic antibody development, while PD-L1 is a critical immune checkpoint widely pursued in cancer immunotherapy—as well as their structural diversity, which allows us to assess the generalizability of our model across distinct binding interfaces.
For each target, we generated several thousand minibinders using SeedProteo. The structures passing the in-silico criteria were further clustered using Foldseek to ensure structural diversity, from which 10 representative minibinders were selected per target for experimental characterization. The selected minibinders were recombinantly expressed in E. coli with an N-terminal His-SUMO tag and purified via Ni-NTA affinity chromatography followed by size-exclusion chromatography (detailed protocols in Appendix 10.1). Binding properties were characterized using surface plasmon resonance (SPR) on a Biacore 8K+ instrument, where candidates were first screened at two concentrations and then subjected to kinetic affinity measurements using either single-cycle or multi-cycle protocols (Appendix 10.2). Protein expression and SPR assays for all designed candidates were performed by WuXi AppTec (China).
As shown in Figure 4, SeedProteo achieved hit rates of 70%–80% across both targets, with the best binders reaching picomolar affinity for PD-L1 and sub-nanomolar affinity for SC2RBD. These results demonstrate that our in-silico design pipeline reliably produces binders with strong experimental binding, bridging the gap between computational prediction and wet-lab validation.
@@ 5 Conclusion
In this work, we presented SeedProteo, a diffusion-based all-atom protein design model. Our empirical benchmarks demonstrate that SeedProteo exhibits superior robustness across long sequences and complex topologies in unconditional design. Furthermore, it achieves state-of-the-art performance among open-source methods for binder design, particularly on hard and multi-chain targets. While effective co-design remains an open challenge in the field, we offer specific insights into achieving superior sequence-structure consistency via a global MRF module. Crucially, our wet-lab validation demonstrates that these in-silico advantages carry over to real-world efficacy, yielding experimentally confirmed binders with high hit rates and picomolar-level affinities. We hope these findings will inspire significant community efforts to further advance the frontier of simultaneous sequence-structure generation.
6 Revision History
Version 2, February 24, 2026
  • Added in vitro experimental validation results.
    • Added wet-lab experiments results with SPR binding characterization on SC2RBD and PD-L1 targets.
    • Added revision history section.
  • Updated several citations.
Version 1, December 31, 2025
Initial manuscript release.